OpenStudy (anonymous):

6. Find the solution to x2 = 121. A. ±121 B. -11 C. ±11 can you also explain why it's the correct answer? So far, I have: x2=± √(121) take square root both the sides than what do I do?

OpenStudy (swissgirl):

alrighty as u said we take the sqrt of both sides \( \sqrt{x^2} =± \sqrt{121} \)

OpenStudy (swissgirl):

\(x=± \sqrt{121}\) what is sqrt(121)?

OpenStudy (swissgirl):

x=±11

OpenStudy (swissgirl):

There is ur answer

OpenStudy (swissgirl):

your welcome